OpenStudy (anonymous):

express as a single logarithm 6ln (x-3)- 5ln(x) please show me step by step... thanks:)...

OpenStudy (anonymous):

\[\ln (x-3)^6-\ln x^5=\ln \frac{(x-3)^6}{x^5}\]

OpenStudy (anonymous):

oh thanks... i got the same answer too:)..

OpenStudy (anonymous):

:)